On average 3 accidents happens every week in a certain city. What is the probability that we will observe 9 accidents in a month


1
Expert's answer
2021-10-29T02:53:02-0400

X ~ Poisson( λ=3)

"P(X=x)= \frac{e^{- λ} λ^x}{x!} \\

P(X=9) = \frac{e^{-3} 3^9}{9!} \\

= \frac{0.04978 \times 19683}{362880} \\

= 0.0027"